OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (anonymous):

A. 2 B. 3/4 C. -4/3 D. -2

OpenStudy (alexandervonhumboldt2):

use the slope formula: \[\frac{ y_{1}-y_{2} }{ x_{1}-x_{2} }\]

OpenStudy (alexandervonhumboldt2):

substitute the values: \[\frac{ 3+7 }{ -1-4 }\]

OpenStudy (alexandervonhumboldt2):

10/(-5)=-2

OpenStudy (anonymous):

so 3 - (-7) divied by (-1) - 4?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

-2 is what I got